A company had sales of $5,250,000 (500,000 units); Variable costs ratio of .65 and Fixed costs of $1,500,000....

-Variable costs can be reduced by $1 per unit if the company incurs an additional $200,000 in fixed costs per year. What will be the effect on income if the company makes the change, and everything else stays the same?

-If sales increase by 20,000 units, the variable cost ratio remains the same, and there is no change in fixed costs or the sales price, what will be the effect on income?

-If the variable cost ratio is reduced to .55 and there is no change in sales volume, price, or fixed costs, what will be the effect on income?

-If fixed costs increase by 8%, and there is no change in sales or variable costs, what will be the effect on income?
Answer #1

Answer

Current selling price = 5,250,000 / 500,000 = $10.5

Variable cost per unit = 10.5 × .65 = $6.825

Fixed cost = $1,500,000

Current operating income

Particulars $
Sales 5,250,000
Less variable cost (500,000 × 6.825) 3,412,500
Contribution margin 1,837,500
Less fixed cost 1,500,000
Operating income 337,500

  

1 variable cost = 6.825 - 1 = $5.825

Fixed cost = 1,500,000 + 200,000 = $1,700,000

Calculation of operating income

Particulars $
Sales 5,250,000
Less variable cost (500,000 × 5.825) 2,912,500
Contribution margin 2,337,500
Less fixed cost 1,700,000
Operating income 637,500

Operating income increase from $337,500 to to 637,500.

2 new selling units = 500,000 - 20,000 = 480,000 units

Calculation of operating income

Particulars $
Sales (480,000 × 10.5) 5,040,000
Less variable cost (480,000 × 6.825) 3,276,000
Contribution margin 1,764,000
Less fixed cost 1,500,000
Operating income 264,000

Operating income decrease from $337,500 to $264,000.

3 variable cost = 10.5 × .55 = $5.775

Calculation of operating income

Particulars $

Sales

5,250,000
Less variable cost (500,000 × 5.775) 2,887,500
Contribution margin 2,362,500
Less fixed cost 1,500,000
Operating income 862,500

4 Fixed cost = 1,500,000 + 8% = $1,620,000

Calculation of operating income

Particulars $
Sales 5,250,000
Less variable cost (500,000 × 6.825) 3,412,500
Contribution margin 1,837,500
Less fixed cost 1,620,000
Operating income 217,0500

Operating income decrease from $337,500 to $217,500.

Effects on operating income on the basis of above changes are calculated above.

If YOU are like the answer please upvote.